# From Time-inconsistency to Time-consistency for Optimal Stopping Problems thanks: The first author would like to acknowledge the funding of National Science Foundation of China (Grant No. 12271462) and Shenzhen Science and Technology Program (2022 College Stable Support Project)

### v1

@@ Abstract
For optimal stopping problems with time-inconsistent preference, we measure the inherent level of time-inconsistency by taking the time needed to turn the naive strategies into the sophisticated ones. In particular, when in a repeated experiment the naive agent can observe her actual sequence of actions which are inconsistent with what she has planned at the initial time, she then chooses her immediate action based on the observations on her later actual behavior. The procedure is repeated until her actual sequence of actions are consistent with her plan at any time. We show that for the preference value of cumulative prospect theory, in which the time-inconsistency is due to the probability distortion, the higher the degree of probability distortion, the more severe the level of time-inconsistency, and the more time required to turn the naive strategies into the sophisticated ones.
@@ 1 Introduction
Optimal stopping problems arise in many economic and financial decision-making scenarios – for example, when an entrepreneur completes a project, an investor sells stock, or a gambler quits gambling at the casino. The stopping decision depends on the preference of the entrepreneur, the investor, or the gambler. In a dynamic decision process, the agent plans a sequence of actions to be taken at each time point. Then, it is possible that later when the agent revisits the same optimal stopping problem at a different time than the initial time, she finds that her previously planned sequence of actions are no longer optimal according to her current preference. This phenomenon is called time-inconsistency.
@@
Time-inconsistency can be observed in many dynamic decision problems. In the casino gambling problem, for example, a gambler may first plan her strategy as follows: She would continue gambling if she gains and stop gambling if she loses – termed the “loss-exit” strategy. When she does gambles, her actual behavior may be opposite to what she has planned: She stops gambling if she gains and continues gambling if she loses – termed the “gain-exit” strategy. This is because her preference characterized by cumulative prospect theory leads to time-inconsistency. See Barberis (2012) and Hu et al. (2023) for more detailed discussions. Other possible scenarios that induce time-inconsistency include mean-variance preference, state-dependent reference, or non-exponential discount factors.
@@
Björk and Murgoci (2010) concludes that there are three approaches for dealing with time-inconsistency. First, the agent ignores the time-inconsistency issue and derives the optimal strategy at the initial time to be her strategy. This is called the pre-committed type. Second, the agent constructs a strategy such that her current decision is the best based on expectations on her future decisions. This is called the sophisticated type. Third, the agent continuously derives her current best strategy and deviates from what she has planned before. This is called the naive type.
@@
Huang et al. (2020) first prove that in a one-dimensional diffusion process when the payoff functional satisfies regularity conditions, the sophisticated equilibrium of the stopping problem can be obtained as a fixed point of an operator, which represents strategic reasoning that takes the future selves’ behaviors into account. In other words, the strategic reasoning may turn a naive agent into a sophisticated one.
@@
In this work we consider the time-inconsistent optimal stopping problem with randomization in a discrete-time setting. In a finite time horizon, the sophisticated strategies can be backward derived from the terminal time to the initial time, whereas naive strategies are derived by optimizing the stopping problem at each node. By taking the naive agent’s actual behavior into consideration, the agent’s strategies eventually match with sophisticated strategies after several rounds of training on strategic reasoning. The time required to turn the naive strategies into the sophisticated ones is taken as a measure of time-inconsistency of the problem. In the example of cumulative prospect theory preferences (CPT), the higher the degree of probability distortion, the more severe the level of time-inconsistency, and the more time required to turn the naive strategies into the sophisticated ones.
@@
The rest of this paper is organized as follows. In section 2, we introduce the model of the optimal stopping problem with randomization, time-inconsistent preferences, and the different types of agents. In section 3, we establish iterations to turn the naive strategies into the sophisticated ones as a measure of time-inconsistency. We illustrate the iteration procedure in section 4, using CPT as an example of time-inconsistent preferences. Section 5 provides the conclusion.
@@ 2.1 Optimal stopping
Consider an optimal stopping problem faced by an agent in a discrete-time simple symmetric random walk. Let $\Delta_{0,0}$ stand for the set of all feasible time-state pairs in the simple symmetric random walk up to time $T$. Consider Markovian stopping strategies with external randomization allowed. At time 0 with initial state 0, the agent determines her stopping strategy by choosing a sequence of actions $a_{0,0}=\{a_{0,0}(t,x)\}_{(t,x)\in\Delta_{0,0}}$, where $a_{0,0}(t,x)\in[0,1]$ stands for the probability to stop at node $(t,x)$. If $a_{0}(t,x)=p$ for some $p\in[0,1]$, the agent tosses a (biased) coin to determine whether to stop or not. If the coin lands on heads, the agent chooses to continue; otherwise, she chooses to stop. The probability of tossing a tail is equal to $p$. In particular, if $a_{0,0}(t,x)=1$, the agent chooses to stop for sure at node $(t,x)$; if $a_{0,0}(t,x)=0$, she chooses to continue for sure at node $(t,x)$.
Suppose that the agent tosses a heads at time 0, which means that she would continue. Then, at time 1 and state $j$, where $j\in\{-1,1\}$, let $\Delta_{1,j}$ stand for the set of all the remaining feasible time-state pairs in simple symmetric random walk after time 1 and starting at state $j$, up to time $T$. Suppose that the agent revisits the problem to choose her action, regardless of $a_{0,0}(1,j)$ – the plan she made at time 0 for node $(1,j)$. Again, at time 1 she chooses a sequence of actions denoted by $a_{1,j}=\{a_{1,j}(t,x)\}_{(t,x)\in\Delta,t\geq 1}$, where $a_{1,j}(t,x)\in[0,1]$ stands for the probability to stop at node $(t,x)$ according to the plan made at time 1. This pattern continues until the terminal time $T>0$ if the agent has not stopped yet.
@@ 2.2 Time-inconsistent preferences
Let $V_{t,x}(a)$ denote the preference value of the agent at time $t$ and state $x$ when applying a sequence of actions $a=\{a(t,x)\}$ afterward. The preference value function $V$ is called time-inconsistent if there exist $(t,x)$ and $(t^{\prime},x^{\prime})$ with $t<t^{\prime}$ and $(s,y)\in\Delta_{t,x}\cap\Delta_{t^{\prime},x^{\prime}}$ such that
@@
where $a^{*}_{t,x}$ is the optimal action sequence planned at $(t,x)$:
$\begin{array}[]{cl}a^{*}_{t,x}:=\arg\max_{a\in\mathbb{R}^{|\Delta_{t,x}|}([0,1])}V_{t,x}(a),\end{array}$
where $\mathbb{R}^{n}([0,1])$ stands for the $n$-dimensional vector taking values in $[0,1]$ in a point-wise manner, and $|A|$ is the number of elements in set $A$. In other words, as long as there exist certain differences in the actions between those that are planned at different times, the preference value $V$ is time-inconsistent. It is straightforward to verify that the total number of nodes in the set $\Delta_{t,x}$ is $(T-t+1)(T-t+2)/2$. Note that there exists a one-to-one correspondence between the elements in the $|\Delta_{t,x}|$-dimensional vector $a$ to the action $a(s,y)$ taken at node $(s,y)$ for some value $s\geq t$ and $y\in\{-s,-(s-2),...s-2,s\}$. In particular, with current time $t$ and state $x$, the $j$-th element in the vector $a$ corresponds to $a_{t,x}(s,y)$, where $s$ is such that $(s-t)(s-t+1)/2<j\leq(s-t+1)(s-t+2)/2$ and $y=s-2j+(s-t)(s-t+1)+2$. Hence, we do not differentiate the vector $a$ and the action sequence $\{a(t,x)\}$ in the following.
@@ 2.3.1 Naive agent
At time 0, the naive agent determines her plan at that time to be
$\begin{array}[]{cl}a^{*}_{0,0}=\arg\max_{a\in\mathbb{R}^{|\Delta_{0,0}|}([0,1])}V_{0,0}(a).\end{array}$
She then takes action $a^{*}_{0,0}(0,0)$. If she does not stop at time 0 and continues to time 1, then at time 1 with state $j\in\{-1,1\}$, she disregards $a^{*}_{0,0}(1,j)$ – the action she should take at that time according to the optimal plan $a^{*}_{0,0}$ made at time 00. Instead, she determines her strategy at that time to be
$\begin{array}[]{cl}a^{*}_{1,j}=\arg\max_{a\in\mathbb{R}^{|\Delta_{1,j}|}([0,1])}V_{1,j}(a).\end{array}$
@@ 2.3.2 Sophisticated agent
A sophisticated agent realizes the time-inconsistency but has no commitment to any predetermined plan. Hence, the sophisticated agent chooses consistent planning in the sense that she optimizes today by expecting her actions in the future. The agent’s selves at different times are considered to be the players of a game, and a consistent plan becomes an intra-personal equilibrium of the game, from which no selves are willing to deviate. Let $\tilde{a}_{T,X_{T}}$ be the strategy determined by the agent at time $T$ with state $X_{T}$. Note that the agent can only choose to stop at time $T$, that is,
@@
Then, at time $T-1$, the sophisticated agent determines her action taken at time $T-1$ according to the action taken at time $T$:
$\begin{array}[]{cl}\tilde{a}_{T-1,X_{T-1}}=\arg\max_{a\in\mathbb{R}^{|\Delta_{T-1,X_{T-1}}|}([0,1]),a(T,X_{T})=1}V_{T-1,X_{T-1}}(a).\end{array}$
Compared with the decision made by a naive agent at time $T-1$, it is the same because both essentially choose the best action at time $T-1$, which is a single-period problem. The situation is different at time $T-2$. The sophisticated agent determines her action taken at time $T-2$ according to the action taken at time $T-1$ and $T$:
$\begin{array}[]{cl}\tilde{a}_{T-2,X_{T-2}}=\arg\max_{a\in\mathbb{R}^{|\Delta_{T-2,X_{T-2}}|}([0,1]),a(T,X_{T})=1,a(T-1,X_{T-1})=\tilde{a}_{T-1,X_{T-1}}(T-1,X_{T-1})}V_{T-2,X_{T-2}}(a).\end{array}$
@@
The sophisticated agent determines her strategy sequentially until time 0 by
$\begin{array}[]{cl}\tilde{a}_{t,X_{t}}=\arg\max_{a\in\mathbb{R}^{|\Delta_{t,X_{t}}|}([0,1]),a(t+j,X_{t+j})=\tilde{a}_{t+j,X_{t+j}}(t+j,X_{t+j}),j=1,...T-t}V_{t,X_{t}}(a).\end{array}$
@@
$a^{S}=\Big{(}\tilde{a}_{0,0}(0,0),...\tilde{a}_{t,X_{t}}(t,X_{t})...\tilde{a}_{T,X_{T}}(T,X_{T})\Big{)}.$
@@ 2.3.3 Pre-committed agent
$\begin{array}[]{cl}a^{*}_{0,0}=\arg\max_{a\in\mathbb{R}^{|\Delta_{0,0}|}([0,1])}V_{0,0}(a).\end{array}$
@@ 3 Turn to Time-consistency
Suppose that in a repeated experiment, the naive agent can observe her actual stopping behavior, based on which she realizes that her optimal plan made at any time is deviated by her decisions made at future time points. Then at time 0 when she plans a sequence of actions taken in a simple symmetric random walk up to time $T$, she (possibly incorrectly) anticipates that her future selves are going to adopt a certain strategy which is consistent with her observation on her actual stopping strategy. This anticipation changes her optimal solution at time 0 to be as follows:
$\begin{array}[]{cl}a^{(1)}_{0,0}=\arg\max_{a\in\mathbb{R}^{|\Delta_{0,0}|}([0,1]),a(t,X_{t})=a^{N(0)}(t,X_{t}),t=1,...T}V_{0,0}(a).\end{array}$
Compared with the agent without training, there exist additional constraints such that
$a(t,X_{t})=a^{N(0)}(t,X_{t}),\;t=1,...T.$
Suppose that the agent does not stop at time 0 and continues to time 1. Then she revisits the problem by the same logic: she anticipates that her future selves are going to adopt the actual stopping strategy and then chooses her optimal solution at time 1 with state $j\in\{-1,1\}$ accordingly:
$\begin{array}[]{cl}a^{(1)}_{1,j}=\arg\max_{a\in\mathbb{R}^{|\Delta_{1,j}|}([0,1]),a(t,X_{t})=a^{N(0)}(t,X_{t}),t=2,...T}V_{1,j}(a).\end{array}$
At time $t$ with state $X_{t}$, she plans her action to be taken according to her actual stopping strategy afterwards:
$\begin{array}[]{cl}a^{(1)}_{t,X_{t}}=\arg\max_{a\in\mathbb{R}^{|\Delta_{t,X_{t}}|}([0,1]),a(t+j,X_{t+j})=a^{N(0)}(t+j,X_{t+j}),t=1,...T-t}V_{t,X_{t}}(a).\end{array}$
@@
$a^{N(1)}=\Big{(}a^{(1)}_{0,0}(0,0),...a^{(1)}_{t,X_{t}}(t,X_{t}),...a^{(1)}_{T,X_{T}}(T,X_{T})\Big{)}.$
If $a^{N(1)}=a^{S}$, then the naive agent has been successfully trained into a sophisticated one with consistent strategies. If $a^{N(1)}\neq a^{S}$, then the naive agent is still not fully sophisticated and requires more rounds of training. In other words, suppose that after $k$ rounds of training, the naive agent’s actual strategy does not equal the sophisticated agent’s strategy, that is, $a^{N(k)}\neq a^{S}$, $k\geq 1$. Then, at time $t$ with state $X_{t}$, the naive agent at the $(k+1)$-th round plans her action to be taken according to her actual $k$-th round of stopping strategy:
$\begin{array}[]{cl}a^{(k+1)}_{t,X_{t}}=\arg\max_{a\in\mathbb{R}^{|\Delta_{t,X_{t}}|}([0,1]),a(t+j,X_{t+j})=a^{N(k)}(t+j,X_{t+j}),t=1,...T-t}V_{t,X_{t}}(a).\end{array}$
The additional constraints $a(t+j,X_{t+j})=a^{N(k)}(t+j,X_{t+j})$, $t=1,...T-t$, show that the agent anticipates that after time $t$, she is going to behave according to $k$-th stopping strategy.
@@
Consider a $T$-period binomial tree. The naive strategies are the same as the sophisticated ones after $T-1$ rounds of training, that is, $a^{N(T-1)}=a^{S}$.
@@
12:        optimize $V_{t,x}(a)$, subject to $a(s,y)=a^{S}(s,y)$, $s\in[t+1,T]\cap\mathbb{Z}$, $y=s-2*(b-1)$ for $b\in[1,s+1]\cap\mathbb{Z}$;
@@
21:           optimize $V_{t,x}(a)$, subject to $a(s,y)=a^{N(k-1)}(s,y)$, $s\in[t+1,T]\mathbb{Z}$, $y=s-2*(b-1)$ for $b\in[1,s+1]\cap\mathbb{Z}$;

### v2

@@ Abstract
For optimal stopping problems with time-inconsistent preference, we measure the inherent level of time-inconsistency by taking the time required to turn the naive strategies into the sophisticated ones. In particular, when in a repeated experiment, the naive agent can observe her actual sequence of actions that is inconsistent with what she has planned at the initial time, and she would then choose her immediate action based on the observations of her later actual behavior. The procedure is repeated until her actual sequence of actions is consistent with her plan at any time. We show that for the preference value of cumulative prospect theory, in which the time-inconsistency is due to the probability distortion, the higher the degree of probability distortion, the more severe the level of time-inconsistency, and the more time required to turn the naive strategies into the sophisticated ones.
@@ 1 Introduction
Optimal stopping problems arise in many economic and financial decision-making scenarios – for example, when an entrepreneur completes a project, an investor sells stock, or a gambler quits gambling at the casino. The stopping decision depends on the preference of the entrepreneur, the investor, or the gambler. In a dynamic decision process, the agent plans a sequence of actions to be taken at each time point. Then, it is possible that later when the agent revisits the same optimal stopping problem at a different time than the initial time, she finds that her previously planned sequence of actions is no longer optimal according to her current preference. This phenomenon is called time-inconsistency.
@@
Time-inconsistency can be observed in many dynamic decision problems. In the casino gambling problem, for example, a gambler may first plan her strategy as follows: She would continue gambling if she gains and stop gambling if she loses – termed the “loss-exit” strategy. When she does gambles, her actual behavior may be opposite to what she has planned: She stops gambling if she gains and continues gambling if she loses – termed the “gain-exit” strategy. This is because her preference characterized by cumulative prospect theory leads to time-inconsistency. See Barberis (2012), He et al. (2017), He et al. (2019) and Hu et al. (2023) for more detailed discussions. Other possible scenarios that induce time-inconsistency include mean-variance preference, state-dependent reference, or non-exponential discount factors.
@@
Björk and Murgoci (2010) concludes that there are three approaches for dealing with time-inconsistency. First, the agent ignores the time-inconsistency issue and derives the optimal strategy at the initial time to be her strategy. This is called the pre-committed type. Second, the agent constructs a strategy such that her current decision is the best based on expectations on her future decisions. This is called the sophisticated type. Third, the agent continuously derives her current best strategy and deviates from what she has planned before. This is called the naive type. The classification of three types of agents parallels the one used in the literature on hyperbolic discounting, which is one reason for the time-inconsistency; see Machina (1989) for more discussions on three types of agents.
@@
A large literature studies the behavior of three types of agents in various decision problems, but the connections between them and even possible transformations into each other remain unclear. Huang et al. (2020) first prove that in a one-dimensional diffusion process when the payoff functional satisfies regularity conditions, the sophisticated equilibrium of the stopping problem can be obtained as a fixed point of an operator, which represents strategic reasoning that takes the future selves’ behaviors into account. In other words, the strategic reasoning may turn a naive agent into a sophisticated one.
@@
In this work, we consider the time-inconsistent optimal stopping problem with randomization in a discrete-time setting. In particular, we propose a measure of time-inconsistency by measuring the gap between the naive strategy and the sophisticated strategy. In a finite time horizon, we use the binomial tree to describe the underlying state process, where each node represents a pair of time and state. The sophisticated strategies can be backward derived from the terminal time to the initial time, whereas naive strategies are derived by optimizing the stopping problem at each node. By taking the naive agent’s actual behavior into consideration, the agent’s strategies eventually match with sophisticated strategies after several rounds of training on strategic reasoning. In the example of cumulative prospect theory preferences (CPT) – one of the well-known non-expected utility theories, the higher the degree of probability distortion, the more severe the level of time-inconsistency, and the more time required to turn the naive strategies into the sophisticated ones; details are shown in Section 4.
According to the time-inconsistency measure, we design an algorithm to transform the naive strategy into the sophisticated one. The algorithm can be applied to any time-inconsistent stopping problem. In addition to the cumulative prospect theory preferences – where the time-inconsistency is due to the probability distortion – we also consider the present-biased preferences – where the time-inconsistency is due to the non-exponential discount factor. We derive analytical results on how many rounds are needed to achieve the sophisticated strategy from the naive one in the optimal stopping problem with immediate cost and with immediate reward. The transformation on the stopping strategies can also be made in other types of time-inconsistent stopping problem.
In addition to the above mentioned literature, our work is also related to the following literature on time-inconsistent optimal stopping problem with cumulative prospect theory: Xu and Zhou (2012), Ebert and Strack (2015), Ebert and Strack (2018), Henderson et al. (2017), Belomestny and Krätschmer (2017), Henderson et al. (2018). The key difference is that these literatures study the optimal stopping problem in continuous time, while our focus is on discrete time. There is also extensive literature on general optimal stopping problem, e.g., Dayanik and Karatzas (2003), Shiryaev (2007), etc. The following literature discusses the general time-inconsistent decision problem: Strotz (1955-1956), Ekeland and Lazrak (2006), Ekeland and Pirvu (2008), Ebert et al. (2020), Tan et al. (2021), Christensen and Lindensjö (2018), Christensen and Lindensjö (2020), Huang and Nguyen-Huu (2018), Huang and Yu (2021), He and Zhou (2022), etc.
@@
The rest of this paper is organized as follows. In section 2, we introduce the model of the optimal stopping problem with randomization, time-inconsistent preferences, and the different types of agents. In section 3, we establish iterations to turn the naive strategies into the sophisticated ones as a measure of time-inconsistency. We illustrate the iteration procedure in section 4, using CPT as an example of time-inconsistent preferences. Section 5 presents the analytical results of transforming the time-inconsistent strategies into time-consistent ones under the present-biased preferences. Section 6 provides the conclusion.
@@ 2.1 Optimal stopping
Consider an optimal stopping problem faced by an agent in a discrete-time simple symmetric random walk. Let $\Delta_{0,0}$ stand for the set of all feasible time-state pairs in the simple symmetric random walk up to time $T$. Consider Markovian stopping strategies with external randomization allowed. At time 0 with initial state 0, the agent determines her stopping strategy by choosing a sequence of actions ${\bf a}_{0,0}=\{a_{0,0}(t,x)\}_{(t,x)\in\Delta_{0,0}}$, where $a_{0,0}(t,x)\in[0,1]$ stands for the probability to stop at node $(t,x)$. If $a_{0,0}(t,x)=p$ for some $p\in[0,1]$, the agent tosses a (biased) coin to determine whether to stop or not. If the coin lands on heads, the agent chooses to continue; otherwise, she chooses to stop. The probability of tossing a tail is equal to $p$. In particular, if $a_{0,0}(t,x)=1$, the agent chooses to stop for sure at node $(t,x)$; if $a_{0,0}(t,x)=0$, she chooses to continue for sure at node $(t,x)$. See Figure 1 for illustration.
Figure 1: A five-period binomial tree. The pair above each node stands for the current time and state.
Suppose that the agent tosses a head at time 0, which means that she would continue. Then, at time 1 and state $j$, where $j\in\{-1,1\}$, let $\Delta_{1,j}$ stand for the set of all the remaining feasible time-state pairs in simple symmetric random walk after time 1 and starting at state $j$, up to time $T$. Suppose that the agent revisits the problem to choose her action, regardless of $a_{0,0}(1,j)$ – the plan she made at time 0 for node $(1,j)$. Again, at time 1 she chooses a sequence of actions denoted by ${\bf a}_{1,j}=\{a_{1,j}(t,x)\}_{(t,x)\in\Delta,t\geq 1}$, where $a_{1,j}(t,x)\in[0,1]$ stands for the probability to stop at node $(t,x)$ according to the plan made at time 1. This pattern continues until the terminal time $T>0$ if the agent has not stopped yet.
@@ 2.2 Time-inconsistent preferences
Let $V_{t,x}({\bf a})$ denote the preference value of the agent at time $t$ and state $x$ when applying a sequence of actions ${\bf a}=\{a(t,x)\}$ afterward. The preference value function $V$ is called time-inconsistent if there exist $(t,x)$ and $(t^{\prime},x^{\prime})$ with $t<t^{\prime}$ and $(s,y)\in\Delta_{t,x}\cap\Delta_{t^{\prime},x^{\prime}}$ such that
@@
where ${\bf a}^{*}_{t,x}=\{a^{*}_{t,x}(s,y)\}$ is the optimal action sequence planned at $(t,x)$, that is,
$\begin{array}[]{cl}{\bf a}^{*}_{t,x}:=\arg\max_{{\bf a}\in\mathbb{R}^{|\Delta_{t,x}|}([0,1])}V_{t,x}({\bf a}),\end{array}$
where $\mathbb{R}^{n}([0,1])$ stands for the $n$-dimensional vector taking values in $[0,1]$ in a point-wise manner, and $|A|$ is the number of elements in set $A$. In other words, as long as there exist certain differences in the actions between those that are planned at different times, the preference value $V$ is time-inconsistent. It is straightforward to verify that the total number of nodes in the set $\Delta_{t,x}$ is $(T-t+1)(T-t+2)/2$. Note that there exists a one-to-one correspondence between the elements in the $|\Delta_{t,x}|$-dimensional vector ${\bf a}$ to the action $a(s,y)$ taken at node $(s,y)$ for some value $s\geq t$ and $y\in\{-s,-(s-2),...s-2,s\}$. In particular, with current time $t$ and state $x$, the $j$-th element in the vector ${\bf a}$ corresponds to $a_{t,x}(s,y)$, where $s$ is such that $(s-t)(s-t+1)/2<j\leq(s-t+1)(s-t+2)/2$ and $y=s-2j+(s-t)(s-t+1)+2$. Hence, we do not differentiate the vector ${\bf a}$ and the action sequence $\{a(t,x)\}$ in the following.
@@ 2.3.1 Naive agent
At time 0, the naive agent determines her plan at that time to be ${\bf a}^{*}_{0,0}$ that solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{0,0}|}([0,1])}&V_{0,0}({\bf a}).\end{array}$
She then takes action $a^{*}_{0,0}(0,0)$. If she does not stop at time 0 and continues to time 1, then at time 1 with state $j\in\{-1,1\}$, she disregards $a^{*}_{0,0}(1,j)$ – the action she should take at that time according to the optimal plan ${\bf a}^{*}_{0,0}$ made at time 00. Instead, she determines her strategy at that time to be ${\bf a}^{*}_{1,j}$ that solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{1,j}|}([0,1])}&V_{1,j}({\bf a}).\end{array}$
@@ 2.3.2 Sophisticated agent
A sophisticated agent realizes the time-inconsistency but has no commitment to any predetermined plan. Hence, the sophisticated agent chooses consistent planning in the sense that she optimizes today by expecting her actions in the future. The agent’s selves at different times are considered to be the players of a game, and a consistent plan becomes an intra-personal equilibrium of the game, from which no selves are willing to deviate. Let $\tilde{\bf a}_{T,X_{T}}$ be the strategy determined by the agent at time $T$ with state $X_{T}$. Note that the agent can only choose to stop at time $T$, that is,
@@
Then, at time $T-1$, the sophisticated agent determines her action taken at time $T-1$ according to the action taken at time $T$. Denote her plan at time $T-1$ to be $\tilde{\bf a}_{T-1,X_{T-1}}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{T-1,X_{T-1}}|}([0,1])}&V_{T-1,X_{T-1}}({\bf a}),\\ \text{subject to }&a(T,X_{T})=\tilde{a}_{T,X_{T}}(T,X_{T}).\end{array}$
Compared with the decision made by a naive agent at time $T-1$, it is the same because both essentially choose the best action at time $T-1$, which is a single-period problem. The situation is different at time $T-2$. The sophisticated agent determines her action taken at time $T-2$ according to the action taken at time $T-1$ and $T$. Denote her plan at time $T-2$ to be $\tilde{\bf a}_{T-2,X_{T-2}}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{T-2,X_{T-2}}|}([0,1])}&V_{T-2,X_{T-2}}({\bf a}),\\ \text{subject to }&a(T,X_{T})=\tilde{a}_{T,X_{T}}(T,X_{T}),\\ &a(T-1,X_{T-1})=\tilde{a}_{T-1,X_{T-1}}(T-1,X_{T-1}).\end{array}$
@@
The sophisticated agent determines her strategy sequentially until time 0 in the same fashion. Denote her plan at time $t$ to be $\tilde{\bf a}_{t,X_{t}}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{t,X_{t}}|}([0,1])}&V_{t,X_{t}}({\bf a}),\\ \text{subject to }&a(s,X_{s})=\tilde{a}_{s,X_{s}}(s,X_{s}),\;s=t+1,...T.\end{array}$
@@
${\bf a}^{S}=\Big{(}\tilde{a}_{0,0}(0,0),...\tilde{a}_{t,X_{t}}(t,X_{t}),...\tilde{a}_{T,X_{T}}(T,X_{T})\Big{)}.$
@@ 2.3.3 Pre-committed agent
$\begin{array}[]{cl}{\bf a}^{*}_{0,0}=\arg\max_{{\bf a}\in\mathbb{R}^{|\Delta_{0,0}|}([0,1])}V_{0,0}({\bf a}).\end{array}$
@@ 3 Turn to Time-consistency
Suppose that in a repeated experiment, the naive agent can observe her actual stopping behavior, based on which she realizes that her optimal plan made at any time is deviated by her decisions made at future time points. Then at time 0 when she plans a sequence of actions taken in a simple symmetric random walk up to time $T$, she (possibly incorrectly) anticipates that her future selves are going to adopt a certain strategy which is consistent with her observation on her actual stopping strategy. This anticipation changes her optimal solution at time 0 to be ${\bf a}^{(1)}_{0,0}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{0,0}|}([0,1])}&V_{0,0}({\bf a}),\\ \text{subject to }&a(s,X_{s})=a^{N(0)}(s,X_{s}),\;s=1,...T.\end{array}$
Note that compared with the agent without training at time 00, there exist additional constraints such that
$a(s,X_{s})=a^{N(0)}(s,X_{s}),\;s=1,...T.$
Suppose that the agent does not stop at time 0 and continues to time 1. Then she revisits the problem by the same logic: she anticipates that her future selves are going to adopt the actual stopping strategy and then chooses her optimal solution at time 1 with state $j\in\{-1,1\}$ accordingly. Denote her plan at time $1$ to be ${\bf a}^{(1)}_{1,j}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{1,j}|}([0,1])}&V_{1,j}({\bf a}),\\ \text{subject to }&a(s,X_{s})=a^{N(0)}(s,X_{s}),\;s=2,...T.\end{array}$
At time $t$ with state $X_{t}$, she plans her action to be taken according to her actual stopping strategy afterwards. Denote her plan at time $t$ to be ${\bf a}^{(1)}_{t,X_{t}}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{t,X_{t}}|}([0,1])}&V_{t,X_{t}}({\bf a}),\\ \text{subject to }&a(s,X_{s})=a^{N(0)}(s,X_{s}),\;s=t+1,...T.\end{array}$
@@
${\bf a}^{N(1)}=\Big{(}a^{(1)}_{0,0}(0,0),...a^{(1)}_{t,X_{t}}(t,X_{t}),...a^{(1)}_{T,X_{T}}(T,X_{T})\Big{)}.$
If ${\bf a}^{N(1)}={\bf a}^{S}$, then the naive agent has been successfully trained into a sophisticated one with consistent strategies. If ${\bf a}^{N(1)}\neq{\bf a}^{S}$, then the naive agent is still not fully sophisticated and requires more rounds of training. In other words, suppose that after $k$ rounds of training, the naive agent’s actual strategy does not equal the sophisticated agent’s strategy, that is, ${\bf a}^{N(k)}\neq{\bf a}^{S}$, $k\geq 1$. Then, at time $t$ with state $X_{t}$, the naive agent at the $(k+1)$-th round plans her action to be taken according to her actual $k$-th round of stopping strategy. Denote her $(k+1)$-th round’s plan at time $t$ to be ${\bf a}^{(k+1)}_{t,X_{t}}$, which solves
$\begin{array}[]{cl}\max_{{\bf a}\in\mathbb{R}^{|\Delta_{t,X_{t}}|}([0,1])}&V_{t,X_{t}}(a),\\ \text{subject to }&a(s,X_{s})=a^{N(k)}(s,X_{s}),\;s=t+1,...T.\end{array}$
The constraints $a(s,X_{s})=a^{N(k)}(s,X_{s})$ show that the agent anticipates that at time $s\geq t+1$, she is going to behave according to $k$-th stopping strategy.
@@
Consider a $T$-period binomial tree. The naive strategies are the same as the sophisticated ones after $T-1$ rounds of training, that is, ${\bf a}^{N(T-1)}={\bf a}^{S}$.
@@
12:        optimize $V_{t,x}({\bf a})$, subject to $a(s,y)=a^{S}(s,y)$, $s\in[t+1,T]\cap\mathbb{Z}$, $y=s-2*(b-1)$ for $b\in[1,s+1]\cap\mathbb{Z}$;
@@
21:           optimize $V_{t,x}({\bf a})$, subject to $a(s,y)=a^{N(k-1)}(s,y)$, $s\in[t+1,T]\cap\mathbb{Z}$, $y=s-2*(b-1)$ for $b\in[1,s+1]\cap\mathbb{Z}$;